### v1

@@ 4 Factorization of factors ${\mathbb{P}}(\omega^{\delta})$
Case (E) By Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$. By the assumption and Lemma 4.4(a) again, for each $\xi<\kappa$ we have $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})=\mathop{\rm cf}\nolimits(\delta_{\xi})\geq\kappa$; thus, by Theorem 4.7, $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$. $\Box$
5 The posets ${\mathbb{P}}(\omega^{\delta})$: Maximal collapse
5.1 Case (A)
Successors
Here we consider a subcase of Case (A) of Theorem 4.10: the posets of the form ${\mathbb{P}}(\omega^{\delta+1})$, where $\delta$ is an ordinal. By Fact 4.3(a) we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1}))={\mathcal{P}}_{\delta+1}:=(P(\omega^{\delta+1})/{\mathcal{I}}_{\omega^{\delta+1}})^{+}$ and, since by Fact 2.12(a) we have $|\omega^{\delta+1}|=|\delta|$, it follows that $|{\mathcal{P}}_{\delta+1}|\leq|P(\omega^{\delta+1})|=2^{|\delta|}$. By Theorem 4.10 the forcing ${\mathcal{P}}_{\delta+1}$ is $\omega_{1}$-closed, thus preserves $\omega_{1}$. Since ${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi$ and the first iteration collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$, it is natural to ask whether it collapses more cardinals (which, of course, depends on $\delta$, ${\mathbb{P}}(\omega^{\delta})$, $\mathop{\rm cc}\nolimits({\mathbb{P}}(\omega^{\delta+1}))$, etc.). Here we show that if ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$ (examples of such $\delta$’s can be constructed using (C), (D), and (E) of Theorem 4.10), then ${\mathfrak{h}}(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))=\omega_{1}$ and, moreover, under some additional conditions, (for example, if ${\mathfrak{h}}=\omega_{1}$, which holds under GCH and in all Cohen extensions, or if $0^{\sharp}$ does not exist and ${\mathfrak{b}}={\mathfrak{d}}$, e.g. if MA is true) the collapse by ${\mathbb{P}}(\omega^{\delta+1})$ is maximal possible: $2^{|\delta|}$ is collapsed to $\omega_{1}$ and, moreover,
@@ 5.1 Case (A)
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|}).$ (36)
@@
Let $0^{\sharp}$ does not exist let ${\mathbb{P}}$ be a separative $\omega$-distributive forcing and $V_{\mathbb{P}}[G]$ a generic extension by ${\mathbb{P}}$. Then
(a) If ${\mathfrak{b}}={\mathfrak{d}}$ and ${\mathfrak{b}}^{V_{\mathbb{P}}[G]}=\omega_{1}$, then $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$;
(b) If $\mathop{\rm cf}\nolimits^{V_{\mathbb{P}}[G]}({\mathfrak{h}})=\omega_{1}$, then $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$.
Proof. (a) Under the assumptions, in $V$ there is a dominating family ${\mathcal{D}}=\{f_{\alpha}:\alpha<{\mathfrak{b}}\}$ in ${}^{\omega}\omega$ well ordered by $\leq^{*}$ (a scale) and in $V_{\mathbb{P}}[G]$ there is an unbounded family $\{g_{\xi}:\xi<\omega_{1}\}$. For $\xi<\omega_{1}$ let $\alpha_{\xi}<{\mathfrak{b}}$, where $g_{\xi}\leq^{*}f_{\alpha_{\xi}}$ Assuming that the set $\{a_{\xi}:\xi<\omega_{1}\}$ is not cofinal in ${\mathfrak{b}}$ there would be $\alpha<{\mathfrak{b}}$ such that for each $\xi<\omega_{1}$ we have $a_{\xi}\leq\alpha$ and, hence, $g_{\xi}\leq^{*}f_{\alpha_{\xi}}\leq^{*}f_{\alpha}$, which is false. Thus the set $X:=\{a_{\xi}:\xi<\omega_{1}\}$ is cofinal in ${\mathfrak{b}}$. Since $0^{\sharp}$ does not exist in $V_{\mathbb{P}}[G]$ too, by Jensen’s Covering Lemma applied in $V_{\mathbb{P}}[G]$ there is a set $C\in L^{V_{\mathbb{P}}[G]}=L\subset V$ such that $X\subset C$ and $|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$. Clearly, the set $C\cap{\mathfrak{b}}$ is unbounded in ${\mathfrak{b}}$ and belongs to $V$ and, since ${\mathfrak{b}}$ is a regular cardinal (in $V$) we have $|C\cap{\mathfrak{b}}|^{V}={\mathfrak{b}}$. Hence we have $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}=|C\cap{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}\leq|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$ and, by the $\omega$-distributivity of ${\mathbb{P}}$, $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}\geq\omega_{1}$, which gives $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$.
(b) Let $\mathop{\rm cf}\nolimits^{V_{\mathbb{P}}[G]}({\mathfrak{h}})=\omega_{1}$ and, in $V_{\mathbb{P}}[G]$, let $X:=\{a_{\xi}:\xi<\omega_{1}\}$ be a cofinal subset of ${\mathfrak{h}}$. Then, as above, there is a set $C\in L\subset V$ such that $X\subset C$ and $|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$. Again, the set $C\cap{\mathfrak{h}}\in V$ is unbounded in ${\mathfrak{h}}$ and, since ${\mathfrak{h}}$ is a regular cardinal (in $V$) we have $|C\cap{\mathfrak{h}}|^{V}={\mathfrak{h}}$. Hence we have $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}=|C\cap{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}\leq|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$ and, by the $\omega$-distributivity of ${\mathbb{P}}$, $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}\geq\omega_{1}$, which gives $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$. $\Box$
@@
If $\delta\geq\omega_{1}$ is an ordinal and forcing by ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$, then we have
(a) $1_{{\mathbb{P}}(\omega^{\delta+1})}\Vdash|((2^{|\delta|})^{V})\check{\;}|=|({\mathfrak{h}}^{V})\check{\;}|$;

### v2

@@ 3 Factorization of ${\mathbb{P}}(\alpha)$
(E) By Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$. By the assumption and Lemma 4.4(a) again, for each $\xi<\kappa$ we have $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})=\mathop{\rm cf}\nolimits(\delta_{\xi})\geq\kappa$; thus, by Theorem 4.7, ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$. $\Box$
The partial orders ${\mathbb{P}}(\alpha)$. ZFC and consistency results
Theorem 4.11
[…]
(a) If $\delta_{i}$ satisfies (A) or (B), for each $i\leq n$, then the partial order $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is $\sigma$-closed and completely embeds ${{\mathcal{P}}_{\omega}}^{k}$, where $k={\sum_{i=1}^{n}s_{i}}$;
(b) Otherwise, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ completely embeds ${\mathcal{P}}_{\lambda}$, for some uncountable regular $\lambda$ and, hence, collapses at least $\omega_{2}$ to $\omega$. If, in addition, $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})=(2^{|\alpha|})^{+}$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\alpha|})$.
Proof. Let $\langle\delta^{\prime}_{j}:1\leq j\leq k\rangle=\langle\delta_{n},\dots,\delta_{n},\dots,\delta_{1},\dots,\delta_{1}\rangle$, where $\delta_{i}$ appears $s_{i}$ many times, for $i\leq n$. Then $\alpha\cong\sum_{1\leq j\leq k}L_{j}:=L$, where $L_{j}\cong\omega^{\delta^{\prime}_{j}}$, for $j\leq k$, and by Theorem 3.2 we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$.
(a) Under the assumptions, by Theorem 4.10 for each $j\leq k$ the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ is $\sigma$-closed and ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$. By Fact 2.1(b) the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed and by Fact 2.3 ${{\mathcal{P}}_{\omega}}^{k}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$.
(b) Otherwise, by Lemma 4.1 there are a regular $\lambda\geq\omega_{1}$ and $j\leq k$, such that ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta^{\prime}_{j}}))$ and, since $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta^{\prime}_{j}}))\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ we have ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$. If $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})=(2^{|\alpha|})^{+}$, then by Fact 2.6(e) forcing by $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ collapses $2^{|\alpha|}$ to $\omega$ and, hence, $|\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)|=2^{|\alpha|}$. So, by Fact 2.5, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\alpha|})$. $\Box$
Theorem 4.12
If ${\mathfrak{h}}<{\mathfrak{c}}=\omega_{2}=2^{\omega_{1}}$, then $\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$ and
(a) For each non-zero ordinal $\delta<{\mathfrak{c}}$ we have
@@
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\left\{\begin{array}[]{cl}\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}}),&\mbox{ if (A) or (B) holds or }\delta<\omega_{1},\\[2.84526pt] \mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}}),&\mbox{ if (D) or (E) holds;}\end{array}\right.$
(b) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$ or $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}})$, for $\alpha\in[\omega,{\mathfrak{c}})$.
Proof. First, by Facts 2.6(b) and 2.5(a) we have $\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$.
(a) For $\delta<\omega_{1}$ by Theorem 1.1(b) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$. If $\delta\in[\omega_{1},{\mathfrak{c}})$, then regarding Lemma 4.1 Case (C) impossible. If $\delta$ satisfies (A) or (B), then by Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\omega_{1}$-closed and ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$, so, by Fact 2.6(b), $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses ${\mathfrak{c}}$ to $\omega_{1}$. Since ${\mathfrak{c}}^{<\omega_{1}}={\mathfrak{c}}$, by Fact 2.5 we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$. If (D) or (E) holds for $\delta$, then by Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses ${\mathfrak{c}}$ to $\omega$ and, by Fact 2.5, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}})$.
(b) If $\alpha\in[\omega_{1},{\mathfrak{c}})$ and $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$, then by Fact 2.9(c) we have $|\alpha|=|\delta_{n}|=\omega_{1}$ and by Fact 4.3(b) $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))|=|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta_{n}}))|\leq 2^{\omega_{1}}={\mathfrak{c}}$. By Theorem 4.11, either the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is $\sigma$-closed and collapses ${\mathfrak{c}}$ to $\omega_{1}$, or collapses ${\mathfrak{c}}$ to $\omega$; thus, the statement follows from Fact 2.5. $\Box$
E.g., regarding Theorem 4.12(a) we obtain $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$, whenever $\mathop{\rm cf}\nolimits(\delta)\leq\omega$, but also for $\delta=\omega_{1}\omega+\omega_{1}$. We note that it is consistent that ${\mathfrak{c}}=\omega_{2}=2^{\omega_{1}}$ and ${\mathbb{P}}(\omega^{2})$ preserves cardinals (see [10]).
5 The posets ${\mathbb{P}}(\omega^{\delta})$ and collapsing algebras
Our intention is to understand how the posets ${\mathbb{P}}(\alpha)$ look like. By Theorem 3.2 one step in that direction is to regard their separative quotients and to classify the corresponding factors of the form $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))=(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$ up to isomorphism of their Boolean completions. By Theorem 1.1 it is consistent that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ for each non-zero ordinal $\delta<\omega_{1}$ and Theorem 4.12(a) gives an extension for uncountable ordinals.
Collapse to $\omega$
Here we detect some classes of ordinals $\delta$ such that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$. More generally, we have
@@ 5 The posets ${\mathbb{P}}(\omega^{\delta})$ and collapsing algebras
If $\lambda\geq\omega$ is a regular cardinal, $\delta\geq\omega_{1}$ an ordinal, $2^{|\delta|}>\lambda$, and the forcing $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\lambda$-closed and collapses $2^{|\delta|}$ to $\lambda$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$.
Proof. Defining ${\mathbb{B}}:=P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$, by Fact 4.3 we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathbb{B}}^{+}$ and $|{\mathbb{B}}^{+}|\leq 2^{|\delta|}$. Since ${\mathbb{B}}^{+}$ collapses $2^{|\delta|}>\lambda$ to $\lambda$ we have $|{\mathbb{B}}^{+}|=2^{|\delta|}$ and $(2^{|\delta|})^{<\lambda}=2^{|\delta|}$; so, by Fact 2.5, $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$. Clearly, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathbb{B}}^{+}$ implies that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))=\mathop{\rm ro}\nolimits({\mathbb{B}})$. $\Box$